OpenStudy (anonymous):

1. Consider the following difference equation: yt = 0.8yt-1 + εt -0.5εt-1 since the subscripts do not appear in here I will rewrite this equation in words where t = "today" and t-1 = "yesterday". y"today" = 0.8y"yesterday" + epsilon"today" - 0.5epsilon"yesterday". My assignment is to derive the homogenous and particular solutions while showing my work. Is anyone able to help with deriving the particular solution?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (mathteacher1729):

Click 'insert equation' to insert equations. :) Makes things a bit easier to read. :)

OpenStudy (amistre64):

y{t} = 0.8 y{t-1} + ε{t} -0.5ε{t-1} is epsilon a function or a standin?

OpenStudy (amistre64):

the epsilons seem to be part of another reurrsion

OpenStudy (anonymous):

I have attached a document with the assignment. I need to use the method of undetermined coefficients to solve for my particular solution..

OpenStudy (amistre64):

well, if i recall correctly, the homogenous part is: y{t} -.8y{t-1} = 0 right?

OpenStudy (amistre64):

y{t} = .8^n y{0}

OpenStudy (amistre64):

does that seem right?

OpenStudy (anonymous):

I believe that the homogenous piece is y{t} = A(.8^t) where A=some arbitrary constant.

OpenStudy (anonymous):

I believe that the homogenous piece is y{t} = A(.8)^t where A=some arbitrary constant

OpenStudy (amistre64):

yeah, im used to an 'n' :) forgot to adapt for a t

OpenStudy (amistre64):

the rest if it im not to sure about tho..... its been awhile

OpenStudy (anonymous):

Thanks for attempt. I'll keep studying..